OpenStudy (anonymous):

A car is traveled at 50 miles per hour for 150 miles and then traveled at 75 miles per hour for 150 miles. what was the cars average speed?

OpenStudy (dumbcow):

avg speed = total distance/total time total distance = 150 + 150 = 300 total time = 150/50 + 150/75 = 3 + 2 = 5 avg speed = 300/5 = 60